# Congestion-Approximators from the Bottom Up

### v2

@@ 1 Introduction
The famous max-flow min-cut theorem [^1] For the sake of this introduction, a flow problem on a graph is a set of excesses and deficits on vertices and a solution is a set of paths connecting excess and deficits where no edge is used in more than one path. implies that any set of excesses or deficits in a graph can be connected by disjoint paths if and only if every cut in the graph has more capacity in the edges that cross the cut than the “demand” that is required to cross it. One direction is easy to see (if a cut does not have enough capacity, clearly one cannot route demand paths across it), and the other can be established by linear programming duality (or explicitly using an algorithm as was done by Ford and Fulkerson in 1956 [12].)
@@
More generally, the set of feasible demands in a graph is captured by the (exponentially sized) set of all cuts in the graph. Remarkably, it was shown that polynomially many cuts could approximately (within logarithmic factors) characterize the congestion of any set of demands [31, 32]. Indeed, the total size of cuts was nearly-linear [^2] Nearly-linear means $O(m\log^{c}n)$ where $c$ is a constant. Almost-linear means $O(m^{1+\epsilon})$ for some $\epsilon=o(1)$. in [31, 33] and could be computed in nearly-linear time as announced in 2014 by Peng [29].
@@
In this paper, we provide an $\alpha$-congestion-approximator (along with an routing scheme) that can be constructed in nearly-linear time with an $\alpha$ of $O(\log^{10}n)$. The number of logarithmic factors is not optimized but the structure is significantly simpler than the construction by Peng [29].
@@
The top-down frame of the decomposition of [33] requires that one compute approximate maximum flows at the very top level which entailed a non-linear runtime. A breakthrough result of Sherman [36] showed how to use congestion-approximators to find approximate maximum flows. [^3] Simultaneously, Kelner et. al. [17] used a dual version of congestion approximators called oblivous routers to give efficient algorithms for approximate maximum flow. Peng [30] used the result of Sherman [36] inside the top-down partitioning method based on [33] for constructing congestion-approximators. There is a chicken and egg problem here as the two methods need each other, and a costly recursion was used to combine them based on ultra-sparsifiers which we discuss a bit more below.
@@
Our result proceeds in a bottom-up iterative fashion; indeed, the bottom level is a “weak expander decomposition” which can be computed using trivial congestion-approximators consisting of singleton vertices. [^4] We note expander decompositions themselves have found wide application, including in the breakthrough near linear time algorithm of [8] for maximum flow. Also, the high-level idea of our congestion-approximator is similar to hierachical expander decomposition [13], and our techniques may be useful in getting from almost linear to near linear. Then we proceed level by level, using the lower level clusterings as “pseudo”-congestion-approximators for the next. The frame is simple, but “around the edges” both literally (the edges of the clusters) and metaphorically there are technical issues which require some attention.
@@
Still, given the power that congestion-approximators provide with respect to understanding the structure of graphs, finding more efficient constructions is important and an effective bottom-up or clustering approach seems a natural path to follow. We make the first and a substantive step on this path in the decade since the announcement of Peng’s also polylogarithmic nearly linear time algorithm [29].
@@
Recent progress on maximum flow can be compared to developments in nearly-linear time Laplacian solvers [38], which were initially very complex with many logarithmic factors. But over the years, new tools were developed that both improved the running times and allowed for better understanding of graphs as well as having broader application. See, for example, [21, 22, 10, 16]. Progress has been much slower for maximum flow in comparison. Perhaps one reason is that solutions to Laplacian linear systems are $\ell_{2}$ optimizing flows that involve the Euclidean norm which is a fair bit simpler than the $\ell_{\infty}$ norm central to maximum flow or norms based on more general convex bodies. We believe this result to be a step in the process of understanding maximum flows better.
@@ 1.1 Previous work
We note that this theorem was extended to multi-commodity flow in an approximate sense in [24], [20] and [26] where flows and corresponding cuts are shown to be related by an $O(\log n)$ factor. The first paper gives a method for finding the sparsity or conductance of a graph to within an $O(\log n)$ factor by using multicommodity flow to embed a complete graph, and the others extend the techniques to give approximations between cuts and solutions for arbitrary multicommodity flow instances. In a parallel thread, the mixing times of random walks or eigenvalues (which involve flow problems with an $\ell_{2}$ objective) are related to cuts via classical and tremendously impactful results of Cheeger [7]. Combining eigenvalues and linear programming methods through semidefinite programming yields improved approximate relationships between embeddings and cuts of roughly $O(\sqrt{\log n})$ [5, 4]. Fast versions of these methods involve the cut-matching game developed in [2, 18, 28, 35] and used directly in this work.
@@
As we noted previously, one can view the (exponentially sized) set of feasible sets of demands in a graph as being a very general measure of a graph’s capabilities. And as also mentioned above, Räcke in [31] showed that a polynomial sized set of cuts and corresponding pre-computed routings approximately model the congestion required for any of (exponentially many) sets of possible demands.
@@
A bit more formally, [31] provides an oblivious routing scheme, where the scheme can obliviously route any set of demands with no more than $\alpha$ times as much congestion as the optimal routing. Here, oblivious roughly means that for any demand pair the routing is done without considering any other demand pair. [31] also provides a decomposition where for any set of demands the maximum congestion (i.e., ratio of demand crossing to edge capacity) on any cut in the decomposition is within a factor of $\alpha$ of the congestion needed to route those demands.
@@
The value for $\alpha$ in [31] is $O(\log^{3}n)$ but was non-constructive: however, constructive and improved schemes were quickly developed with $\alpha$ of $O(\log^{2}n\log\log n)$ in [6, 15].
@@
An alternative scheme also by Räcke [32], gave a remarkably simple oblivious routing scheme, that consisted of $O(m)$ trees. The oblivious routing was simply to route any demand by splitting the flow among the trees. Of course, each tree implicitly corresponds to a laminar family of cuts. Still the total size of both the routings and the implicit total size of the cuts is quadratic or worse. Moreover, the time complexity for its construction was also at least quadratic. The approximation factor $\alpha$ for this scheme was $O(\log n)$. We will refer to this as Räcke’s tree scheme.
@@
Madry, in [27], gave a nearly-linear time algorithm that produced a congestion-approximator which has almost-linear size and running time at the cost of having approximation factor of $O(n^{\epsilon})$ that is based on Räcke tree scheme. A central idea is the use of ultrasparsifiers which were introduced by Spielman and Teng [38] in their breakthrough results on linear time solvers for Laplacian linear systems. An ultrasparsifier is formed by taking a certain kind of spanning tree (called low-stretch), making clusters from small connected components of the tree, and sampling a very small set of non-tree edges between the clusters. Such a graph approximates the cuts to within a small factor in the sense that cuts in the original graph and the new one have approximately the same size. This allows a (complicated) recursion where one use several ultrasparsifiers to approximate the cuts in the graph and recursively produce oblivious routing schemes for each of the ultrasparsifiers. Again, this approach is more efficient in time than Räcke’s tree scheme at the cost of a worse approximation factor. The scheme enabled the approximate solution to a host of problems in almost linear time. This contribution was remarkable.
@@
In sum, Madry [27] produced a $\alpha=O(n^{\epsilon})$-congestion approximator in the almost linear running time of $O(m^{1+\epsilon})$. [^6] The $\epsilon$ is subconstant, roughly $1/\sqrt{\log n}$ which means an $m^{\epsilon}$ factor is larger than any polylogarithmic factor. The obstruction to getting to polylogarithmic overheads in Madry’s scheme is the need to recurse on several ultrasparsifiers to keep the approximation from blowing up during the recursion.
@@
In another exciting development, Sherman [36] used Madry’s [27] structure to produce an almost-linear time algorithm for $1+\epsilon$-approximate undirected maximum flow. This is remarkable in that he reduced Madry’s approximation factor of $O(n^{\epsilon})$ to $(1+\epsilon)$. The approximation factor in Madry’s construction translates into a running time overhead, which results in the almost-linear running time we mentioned. We point out that there is close relationship to the work of Spielman and Teng [38], who used ultrasparsifiers to spectrally approximate a Laplacian system to provide algorithms whose running times depend on the approximation quality while obtaining arbitrarily precise solutions. In that case, the dependence of running time on precision was logarithmic whereas Sherman’s algorithm suffers a $O(1/\epsilon^{2})$ dependence on the error $\epsilon$. To be sure, in solving linear systems the idea of using an “approximate” graph in improving condition numbers was known as preconditioning, but for maximum flow or $\ell_{1}$ optimization this was a striking development.
@@
Sherman’s method (could be seen) to use the multiplicative weights framework (see [3]) to route flow across cuts so that every cut has near zero residual demand across it, in particular, $\delta=\epsilon/\alpha$ fraction of its capacity. The congestion approximator then ensures that the remaining flow can be routed with $\alpha\cdot\delta=\epsilon$ congestion and one can recurse with slightly more than $\epsilon$ extra congestion.
@@
We note that Sherman’s result (with Madry’s construction) bypasses longstanding barriers to faster maximum flow algorithms. Since the work of Dinitz [11] in 1973, algorithms had to pay either for path length or for number of paths. Dinitz himself traded this off to get a $O(m^{3/2})$ flow algorithm [9, 23]. Using linear time solvers to do electrical flow, one could eke out a few more paths simultaneously, but one still was pretty stuck at $O(m^{4/3})$. The congestion-approximator and the multiplicative weights optimization method bypasses these obstructions.
@@
In 2014, Räcke, Shah, and Taubig [33] made progress on Räcke’s decomposition based approach. In particular, they showed how to produce a decomposition with approximation parameter of $O(\log^{4}n)$ using maximum flow computations of total size $O(m\log^{5}n)$. They use the same frame as Räcke [31], but use single commodity flows in the context of the cut-matching game [19, 28] which, as previously mentioned, can replace multicommodity flows in finding sparse cuts and routings.
@@
Note, that a congestion-approximator can be used to compute flows efficiently where flows can now be used to compute congestion-approximators. This suggests recursion, but the construction in [33] (and indeed previous ones) were very much top down. That is, the first level of the decomposition (which itself is a tour de force) falls victim to the fact that typical paths in the flows that find them can be long and somewhat abundant. Thus, one critically needs something like a congestion-approximator right away to even compute the top level of the decomposition.
@@
Still, Peng [30] was able to get to a nearly-linear running time recursive method by using ultrasparsifiers and Sherman’s approximate maximum flow algorithm. He constructs an ultrasparsifier of size $O(m/\log^{c}m)$ with polylogarithmic approximation factor, recursively computes a cut sparsifier for the resulting graph, and then argues that the combination of Sherman’s algorithm with both the ultrasparsifier and the congestion-approximator can be used to compute an approximate maximum flow in $O(m\log^{c}m)$ time. Again, the key here is that Sherman [36] allows him to use polylogarithmic overhead to combat the polylogarithmic approximation in both the ultrasparsifer and in the congestion-approximator of the ultrasparsifier. Still, the recursion is a bear, resulting in an admittedly unoptimized runtime of $O(m\log^{41}n)$.
@@
As noted previously, our method is bottom up from the start and avoids the costly recursion required above. We have not computed the exponent of the logarithmic factors due to, for example, using the fair cuts method of [25] as a black box. Still, the clustering approach is more natural and we expect is a useful frame.
@@ 2 Technical Overview
    For each $i\in[L-1]$, for each cluster $C\in\mathcal{P}_{i+1}$, the inter-cluster edges of $\mathcal{P}_{i}$ internal to $C$, together with the boundary edges of $C$, mix [^7] Informally, a set of edges mixes if there is a low-congestion multi-commodity flow between the set of edges whose demand pairs form an expander. In other words, there is an expander flow (in the [5] sense) between the set of edges. Our formal definition is in the preliminaries and only considers single-commodity flows. in the graph $G$. Moreover, the mixings over all clusters $C\in\mathcal{P}_{i+1}$ have congestion $\alpha$ simultaneously.
@@
Then, their union $\mathcal{C}=\bigcup_{i\in[L]}\mathcal{R}_{\geq i}$ is a congestion-approximator with quality $5L^{2}\alpha\beta$.
@@
Overall, the partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{L}$ can be viewed as hierarchical expander decompositions where each partition $\mathcal{P}_{i+1}$ is a weak-expander decomposition relative to the inter-cluster edges of partition $\mathcal{P}$. It is instructive to compare our partitioning to the expander hierarchy construction of [14], where clusters of the previous partition are contracted before building the next partition. While [14] can also extract a congestion approximator from their expander hierarchy construction, their quality is $n^{o(1)}$ because they lose a multiplicative factor from contracting vertices on each level of the hierarchy. To avoid this multiplicative blow-up per level, we do not contract vertices, so our partitioning is not truly hierarchical: a cluster in partition $\mathcal{P}_{i}$ may be “cut” into many components by the next partition $\mathcal{P}_{i+1}$ (see Figure 1, left). While a hierarchical construction is not required, it is a useful property to have when analyzing the quality of the final congestion-approximator. To establish a hierarchy, our key insight is to consider the common refinement of partitions $\mathcal{P}_{i},\mathcal{P}_{i+1},\ldots,\mathcal{P}_{L}$, which we name $\mathcal{R}_{\geq i}$. The partitions $\mathcal{R}_{\geq i}$ over all $i$ are now hierarchical by construction, and we show that the union of all refinements $\mathcal{R}_{\geq i}$ is a congestion-approximator with good quality.
We emphasize that our conceptual idea of not contracting clusters and looking at common refinements is novel and may have future applications to bottom-up constructions of hierarchical objects, especially for obtaining polylogarithmic approximations, meaning that one cannot lose a multiplicative factor at each level. Given that [14] has popularized bottom-up hierarchical approaches, and that their methods so far can only obtain $n^{o(1)}$-factors due to multiplicative errors across levels, we believe that our ideas are promising for future development in this area.

### v3

@@ 1 Introduction
The famous max-flow min-cut theorem [^1] For the sake of this introduction, a flow problem on a graph is a set of excesses and deficits on vertices and a solution is a set of paths connecting excess and deficits where no edge is used in more than one path. implies that any set of excesses or deficits in a graph can be connected by disjoint paths if and only if every cut in the graph has more capacity in the edges that cross the cut than the “demand” that is required to cross it. One direction is easy to see (if a cut does not have enough capacity, clearly one cannot route demand paths across it), and the other can be established by linear programming duality (or explicitly using an algorithm as was done by Ford and Fulkerson in 1956 [FF56].)
@@
More generally, the set of feasible demands in a graph is captured by the (exponentially sized) set of all cuts in the graph. Remarkably, it was shown that polynomially many cuts could approximately (within logarithmic factors) characterize the congestion of any set of demands [Räc02, Räc08]. Indeed, the total size of cuts was nearly-linear [^2] Nearly-linear means $O(m\log^{c}n)$ where $c$ is a constant. Almost-linear means $O(m^{1+\epsilon})$ for some $\epsilon=o(1)$. in [Räc02, RST14] and could be computed in nearly-linear time as announced in 2014 by Peng [Pen14].
@@
In this paper, we provide an $\alpha$-congestion-approximator (along with an routing scheme) that can be constructed in nearly-linear time with an $\alpha$ of $O(\log^{10}n)$. The number of logarithmic factors is not optimized but the structure is significantly simpler than the construction by Peng [Pen14].
@@
The top-down frame of the decomposition of [RST14] requires that one compute approximate maximum flows at the very top level which entailed a non-linear runtime. A breakthrough result of Sherman [She13] showed how to use congestion-approximators to find approximate maximum flows. [^3] Simultaneously, Kelner et. al. [KLOS14] used a dual version of congestion approximators called oblivous routers to give efficient algorithms for approximate maximum flow. Peng [Pen16] used the result of Sherman [She13] inside the top-down partitioning method based on [RST14] for constructing congestion-approximators. There is a chicken and egg problem here as the two methods need each other, and a costly recursion was used to combine them based on ultra-sparsifiers which we discuss a bit more below.
@@
Our result proceeds in a bottom-up iterative fashion; indeed, the bottom level is a “weak expander decomposition” which can be computed using trivial congestion-approximators consisting of singleton vertices. [^4] We note expander decompositions themselves have found wide application, including in the breakthrough near linear time algorithm of [CKL+22] for maximum flow. Also, the high-level idea of our congestion-approximator is similar to hierachical expander decomposition [GRST21a], and our techniques may be useful in getting from almost linear to near linear. Then we proceed level by level, using the lower level clusterings as “pseudo”-congestion-approximators for the next. The frame is simple, but “around the edges” both literally (the edges of the clusters) and metaphorically there are technical issues which require some attention.
@@
Still, given the power that congestion-approximators provide with respect to understanding the structure of graphs, finding more efficient constructions is important and an effective bottom-up or clustering approach seems a natural path to follow. We make the first and a substantive step on this path in the decade since the announcement of Peng’s also polylogarithmic nearly linear time algorithm [Pen14].
@@
Recent progress on maximum flow can be compared to developments in nearly-linear time Laplacian solvers [ST04], which were initially very complex with many logarithmic factors. But over the years, new tools were developed that both improved the running times and allowed for better understanding of graphs as well as having broader application. See, for example, [KMP11, KS16, CKM+14, JS21]. Progress has been much slower for maximum flow in comparison. Perhaps one reason is that solutions to Laplacian linear systems are $\ell_{2}$ optimizing flows that involve the Euclidean norm which is a fair bit simpler than the $\ell_{\infty}$ norm central to maximum flow or norms based on more general convex bodies. We believe this result to be a step in the process of understanding maximum flows better.
@@ 1.1 Previous work
We note that this theorem was extended to multi-commodity flow in an approximate sense in [LR99], [KRAR95] and [LLR95] where flows and corresponding cuts are shown to be related by an $O(\log n)$ factor. The first paper gives a method for finding the sparsity or conductance of a graph to within an $O(\log n)$ factor by using multicommodity flow to embed a complete graph, and the others extend the techniques to give approximations between cuts and solutions for arbitrary multicommodity flow instances. In a parallel thread, the mixing times of random walks or eigenvalues (which involve flow problems with an $\ell_{2}$ objective) are related to cuts via classical and tremendously impactful results of Cheeger [Che70]. Combining eigenvalues and linear programming methods through semidefinite programming yields improved approximate relationships between embeddings and cuts of roughly $O(\sqrt{\log n})$ [ARV09, ALN05]. Fast versions of these methods involve the cut-matching game developed in [AHK10, KRV09a, OSVV08, She09] and used directly in this work.
@@
As we noted previously, one can view the (exponentially sized) set of feasible sets of demands in a graph as being a very general measure of a graph’s capabilities. And as also mentioned above, Räcke in [Räc02] showed that a polynomial sized set of cuts and corresponding pre-computed routings approximately model the congestion required for any of (exponentially many) sets of possible demands.
@@
A bit more formally, [Räc02] provides an oblivious routing scheme, where the scheme can obliviously route any set of demands with no more than $\alpha$ times as much congestion as the optimal routing. Here, oblivious roughly means that for any demand pair the routing is done without considering any other demand pair. [Räc02] also provides a decomposition where for any set of demands the maximum congestion (i.e., ratio of demand crossing to edge capacity) on any cut in the decomposition is within a factor of $\alpha$ of the congestion needed to route those demands.
@@
The value for $\alpha$ in [Räc02] is $O(\log^{3}n)$ but was non-constructive: however, constructive and improved schemes were quickly developed with $\alpha$ of $O(\log^{2}n\log\log n)$ in [BKR03, HHR03].
@@
An alternative scheme also by Räcke [Räc08], gave a remarkably simple oblivious routing scheme, that consisted of $O(m)$ trees. The oblivious routing was simply to route any demand by splitting the flow among the trees. Of course, each tree implicitly corresponds to a laminar family of cuts. Still the total size of both the routings and the implicit total size of the cuts is quadratic or worse. Moreover, the time complexity for its construction was also at least quadratic. The approximation factor $\alpha$ for this scheme was $O(\log n)$. We will refer to this as Räcke’s tree scheme.
@@
Madry, in [Mad10], gave a nearly-linear time algorithm that produced a congestion-approximator which has almost-linear size and running time at the cost of having approximation factor of $O(n^{\epsilon})$ that is based on Räcke tree scheme. A central idea is the use of ultrasparsifiers which were introduced by Spielman and Teng [ST04] in their breakthrough results on linear time solvers for Laplacian linear systems. An ultrasparsifier is formed by taking a certain kind of spanning tree (called low-stretch), making clusters from small connected components of the tree, and sampling a very small set of non-tree edges between the clusters. Such a graph approximates the cuts to within a small factor in the sense that cuts in the original graph and the new one have approximately the same size. This allows a (complicated) recursion where one use several ultrasparsifiers to approximate the cuts in the graph and recursively produce oblivious routing schemes for each of the ultrasparsifiers. Again, this approach is more efficient in time than Räcke’s tree scheme at the cost of a worse approximation factor. The scheme enabled the approximate solution to a host of problems in almost linear time. This contribution was remarkable.
@@
In sum, Madry [Mad10] produced a $\alpha=O(n^{\epsilon})$-congestion approximator in the almost linear running time of $O(m^{1+\epsilon})$. [^6] The $\epsilon$ is subconstant, roughly $1/\sqrt{\log n}$ which means an $m^{\epsilon}$ factor is larger than any polylogarithmic factor. The obstruction to getting to polylogarithmic overheads in Madry’s scheme is the need to recurse on several ultrasparsifiers to keep the approximation from blowing up during the recursion.
@@
In another exciting development, Sherman [She13] used Madry’s [Mad10] structure to produce an almost-linear time algorithm for $1+\epsilon$-approximate undirected maximum flow. This is remarkable in that he reduced Madry’s approximation factor of $O(n^{\epsilon})$ to $(1+\epsilon)$. The approximation factor in Madry’s construction translates into a running time overhead, which results in the almost-linear running time we mentioned. We point out that there is close relationship to the work of Spielman and Teng [ST04], who used ultrasparsifiers to spectrally approximate a Laplacian system to provide algorithms whose running times depend on the approximation quality while obtaining arbitrarily precise solutions. In that case, the dependence of running time on precision was logarithmic whereas Sherman’s algorithm suffers a $O(1/\epsilon^{2})$ dependence on the error $\epsilon$. To be sure, in solving linear systems the idea of using an “approximate” graph in improving condition numbers was known as preconditioning, but for maximum flow or $\ell_{1}$ optimization this was a striking development.
@@
Sherman’s method (could be seen) to use the multiplicative weights framework (see [AHK12]) to route flow across cuts so that every cut has near zero residual demand across it, in particular, $\delta=\epsilon/\alpha$ fraction of its capacity. The congestion approximator then ensures that the remaining flow can be routed with $\alpha\cdot\delta=\epsilon$ congestion and one can recurse with slightly more than $\epsilon$ extra congestion.
@@
We note that Sherman’s result (with Madry’s construction) bypasses longstanding barriers to faster maximum flow algorithms. Since the work of Dinitz [Din06] in 1973, algorithms had to pay either for path length or for number of paths. Dinitz himself traded this off to get a $O(m^{3/2})$ flow algorithm [CKM+11, LRS13]. Using linear time solvers to do electrical flow, one could eke out a few more paths simultaneously, but one still was pretty stuck at $O(m^{4/3})$. The congestion-approximator and the multiplicative weights optimization method bypasses these obstructions.
@@
In 2014, Räcke, Shah, and Taubig [RST14] made progress on Räcke’s decomposition based approach. In particular, they showed how to produce a decomposition with approximation parameter of $O(\log^{4}n)$ using maximum flow computations of total size $O(m\log^{5}n)$. They use the same frame as Räcke [Räc02], but use single commodity flows in the context of the cut-matching game [KRV09b, OSVV08] which, as previously mentioned, can replace multicommodity flows in finding sparse cuts and routings.
@@
Note, that a congestion-approximator can be used to compute flows efficiently where flows can now be used to compute congestion-approximators. This suggests recursion, but the construction in [RST14] (and indeed previous ones) were very much top down. That is, the first level of the decomposition (which itself is a tour de force) falls victim to the fact that typical paths in the flows that find them can be long and somewhat abundant. Thus, one critically needs something like a congestion-approximator right away to even compute the top level of the decomposition.
@@
Still, Peng [Pen16] was able to get to a nearly-linear running time recursive method by using ultrasparsifiers and Sherman’s approximate maximum flow algorithm. He constructs an ultrasparsifier of size $O(m/\log^{c}m)$ with polylogarithmic approximation factor, recursively computes a cut sparsifier for the resulting graph, and then argues that the combination of Sherman’s algorithm with both the ultrasparsifier and the congestion-approximator can be used to compute an approximate maximum flow in $O(m\log^{c}m)$ time. Again, the key here is that Sherman [She13] allows him to use polylogarithmic overhead to combat the polylogarithmic approximation in both the ultrasparsifer and in the congestion-approximator of the ultrasparsifier. Still, the recursion is a bear, resulting in an admittedly unoptimized runtime of $O(m\log^{41}n)$.
@@
As noted previously, our method is bottom up from the start and avoids the costly recursion required above. We have not computed the exponent of the logarithmic factors due to, for example, using the fair cuts method of [LNPS23] as a black box. Still, the clustering approach is more natural and we expect is a useful frame.
@@ 2 Technical Overview
    For each $i\in[L-1]$, for each cluster $C\in\mathcal{P}_{i+1}$, the inter-cluster edges of $\mathcal{P}_{i}$ internal to $C$, together with the boundary edges of $C$, mix [^7] Informally, a set of edges mixes if there is a low-congestion multi-commodity flow between the set of edges whose demand pairs form an expander. In other words, there is an expander flow (in the [ARV09] sense) between the set of edges. Our formal definition is in the preliminaries and only considers single-commodity flows. in the graph $G$. Moreover, the mixings over all clusters $C\in\mathcal{P}_{i+1}$ have congestion $\alpha$ simultaneously.
@@
Then, their union $\mathcal{C}=\bigcup_{i\in[L]}\mathcal{R}_{\geq i}$ is a congestion-approximator with quality $16L^{2}\alpha\beta$.
@@
Overall, the partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{L}$ can be viewed as hierarchical expander decompositions where each partition $\mathcal{P}_{i+1}$ is a weak-expander decomposition relative to the inter-cluster edges of partition $\mathcal{P}$. It is instructive to compare our partitioning to the expander hierarchy construction of [GRST21b], where clusters of the previous partition are contracted before building the next partition. While [GRST21b] can also extract a congestion approximator from their expander hierarchy construction, their quality is $n^{o(1)}$ because they lose a multiplicative factor from contracting vertices on each level of the hierarchy. To avoid this multiplicative blow-up per level, we do not contract vertices, so our partitioning is not truly hierarchical: a cluster in partition $\mathcal{P}_{i}$ may be “cut” into many components by the next partition $\mathcal{P}_{i+1}$ (see Figure 1, left). While a hierarchical construction is not required, it is a useful property to have when analyzing the quality of the final congestion-approximator. To establish a hierarchy, our key insight is to consider the common refinement of partitions $\mathcal{P}_{i},\mathcal{P}_{i+1},\ldots,\mathcal{P}_{L}$, which we name $\mathcal{R}_{\geq i}$. The partitions $\mathcal{R}_{\geq i}$ over all $i$ are now hierarchical by construction, and we show that the union of all refinements $\mathcal{R}_{\geq i}$ is a congestion-approximator with good quality.
We emphasize that our conceptual idea of not contracting clusters and looking at common refinements is novel and may have future applications to bottom-up constructions of hierarchical objects, especially for obtaining polylogarithmic approximations, meaning that one cannot lose a multiplicative factor at each level. Given that [GRST21b] has popularized bottom-up hierarchical approaches, and that their methods so far can only obtain $n^{o(1)}$-factors due to multiplicative errors across levels, we believe that our ideas are promising for future development in this area.